OpenStudy (anonymous):

A man walks 10 miles in x hours. How far does he walk in one hour

OpenStudy (anonymous):

\[10m-xhrs\] \[?m-1hr\]

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

10 miles in 5 hours, \(\frac{10}{5}=2\) mph 10 miles in 2 hours, \(\frac{10}{2}=5\) mph 10 miles in \(x\) hours \(\frac{10}{x}\) mph

OpenStudy (anonymous):

there is no such thing as "cross multiply"

OpenStudy (anonymous):

the answer is the same if you cross mulltiply, might have a different meaning for you